A particle starting from rest moves with constant acceleration. If it takes 5.0 s to reach the speed 18.0 km/h find (a) the average velocity during this period, and (b) the distance travelled by the particle during this period.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. u = 0, v = 18 km/hr = 5 m/s, t = 5 sec a = (v-u )/t = (5-0)/5 = 1 m/s2. s = ut+ ½ at2 = 12.5 m a) Average velocity Vave = (12.5)/5 = 2.5 m/s b) Distance travelled is 12.5 m.